OpenStudy (anonymous):

im trying to solve P(n,3) = 2(n-1 P 3) using permutations

hartnn (hartnn):

and where are u stuck ?

OpenStudy (anonymous):

how to solve each side to get an answer

hartnn (hartnn):

you know the definition of \(\large _nP^r\) ?

hartnn (hartnn):

sorry, \(\large ^nP_r\)

OpenStudy (anonymous):

i know its permutations

hartnn (hartnn):

so, \(\huge \dfrac{n!}{(n-3)!}=2 \times \dfrac{(n-1)!}{(n-4)!}\) did u get till here ?

OpenStudy (anonymous):

yes, i don't know how to show to solve each side from there

hartnn (hartnn):

use the fact that \(\huge n!=n(n-1)!\) and \(\huge n!=n(n-1)!=n(n-1)(n-2)!\) and so on...

OpenStudy (anonymous):

how do i solve the right side of the right side f the equation when its multiplied by 2?

hartnn (hartnn):

\(\huge \dfrac{n(n-1)!}{(n-3)(n-4)!}=2 \times \dfrac{(n-1)!}{(n-4)!}\) did you get what i did ?

hartnn (hartnn):

numerator of left is obvious, for denominator, \(\large (n-3)! = (n-3)(n-3-1)! = (n-3)(n-4)!\) got this ?

OpenStudy (anonymous):

i dont understand why the denominator is like that

hartnn (hartnn):

\(\large a!=a(a-1)!\) this is standard, right ? now just replace a by n-3

OpenStudy (anonymous):

ok

hartnn (hartnn):

now what gets cancelled ?

OpenStudy (anonymous):

i dont know

hartnn (hartnn):

now you got this right ? \(\huge \dfrac{n(n-1)!}{(n-3)(n-4)!}=2 \times \dfrac{(n-1)!}{(n-4)!}\) how can you simplify this ? notice that some terms are getting cancelled!

OpenStudy (anonymous):

does the (n-1)! on the tops cancel out and the (n-4)! on the bottoms cancell out?

hartnn (hartnn):

YES! and what remains ??

OpenStudy (anonymous):

\[\frac{ n }{ (n-3) } = 2\]

hartnn (hartnn):

yes, isn't this easy to solve ? (NOTE : since we are cancelling (n-1)! , which can be 0, we need to account for n-1 =0 and so, n=1 is a solution)

hartnn (hartnn):

and since we cancelled, (n-4) ! also, solve (n-4)! = 0, n=.... ?, ...?

hartnn (hartnn):

are u getting all this ?

OpenStudy (anonymous):

so becasue it equaled the fraction, then it goes down to 2(n-3)=n?

OpenStudy (anonymous):

and then it goes to 2n-6=n ?

hartnn (hartnn):

yes, continue, n=... ?

OpenStudy (anonymous):

2n=n+6 ?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

or is it n =6

OpenStudy (anonymous):

thanks for your help!

hartnn (hartnn):

n=6 is correct! :) welcome ^_^ and \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)